particle is moving along the line 4y+3x=5 in gravitational field intensity 3i+4j work done in moving particle through 10m on line is

Grade:12

2 Answers

Vikas TU
7 years ago
In x direcn.
dWx = 3*(5-3x)/4dx
On ntegerating we get,
Wx = 0.75(5x – 1.5x^2)
put x = 10m weget,
Wx = -75 Joule
In y direction,
dWy = 4*(5 – 4y)/3dy
On integrating we get,
Wy =  4(5y – 2y^2)/3
put y = 10 we get,
Wy = > -200 J
Sanjeev Srivastava
5 years ago
Gravitational field or E = (3i+4j)
F = mE = 1 * (3i+4j)
W = F. d = (3i+4j) . d  --------- (1)
// For d =>
y = -(3/4)x + 5/4
Tan β = - 3/4 = 37°(x-axis negative) 
Solving the slope for 10m in vector form →
(-10sinβi +10cosβj) => (-8i + 6j) // d.
Putting value back in (1) and solving the dot product. 
W = - 24 + 24 = 0.